# Harnessing Test-time Adaptation for NLU tasks Involving Dialects of English

### v1

@@ 4 Experiments
We also compare the performance of using SHOT versus performing finetuning on dialectal data. Table 3 shows that SHOT always improves the performance of models trained on SAE data. Furthermore, SHOT commonly out-performances model trained on dialectal data. This suggests that building dialect-robust models require the incorporation of SHOT-based self-supervised techniques in order to reverse the effect of performing self-supervised learning of base embedding models on SAE data.
When analyzing the relationship between dialects models are finetuned on and their performance on varying dialects, we also observe the asymmetry of evaluation. Specifically, as shown on Table 2, we find that when training on a dialect $A$, its performance is best when evaluating on dialect $A$ (with the exception of SAE, see italicised entries on Table 2). However, when training on dialect $A$, $B_{1}$, $B_{2}$, etc., and evaluating on dialect $A$, oftentimes, training on a dialect $B_{i}$ yields higher performance than training on dialect $A$. Specifically, SAE shows the best performance with or without test-time adaptation in multiple dialects (see boldened entries in Table 2).
@@ 5 Conclusion
Due to the dialectal gap which exhibits in previous literature, we formally propose a mathematical definition of the concept, apply SHOT Liang et al. (2020) to dialectal NLP, and do a comparative study of SHOT against finetuning under different dialectal settings. Our findings indicate that: (1) There is a strong positive relationship between the dialectal gap and $\Delta$. This means that as the dialectal gap increases, the more TTA will be beneficial in reducing said gap. (2) SHOT always improves the performance of models trained on SAE data. Furthermore, SHOT commonly out-performances model trained on dialectal data. (3) When training on a dialect $A$, its performance is best when evaluating on dialect $A$. However, when training on dialect $A$, $B_{1}$, $B_{2}$, etc., and evaluating on dialect $A$, oftentimes, training on a dialect $B_{i}$ yields higher performance than training on dialect $A$.
@@
We are unable to perform evaluation on other GLUE tasks due to their dataset sizes being significantly higher than our computational requirement. We also weren’t able to do experiments on STS-B due to this being a regression task which is not yet accommodated by SHOT Liang et al. (2020). We use Multi-VALUE to create dialectal transformations. We acknowledge that there may be errors. Similarly, dialects are a combination of syntax, vocabulary, orthography and culture. These nuances are not captured in our evaluative datasets. RoBERTA was pretrained predominantly on SAE corpus, we have not explored the effect of RoBERTA or other language models if pretrained with dialectal data predominantly.
Acknowledgments
The production of the code used in this paper was partially reliant on generative AI code assistant. Generative AI tools were also used to perform literature search. All AI-generation has been validated by the authors. This work was funded by Duke Nguyen’s PhD Scholarship at the University of New South Wales, Australia. Dedicated to the memory of Ryan Herford (1997-2024).

### v2

@@ 3.4 Results
When comparing the performance of using SHOT versus performing finetuning on dialectal data, we find that SHOT always improves the performance of models trained on SAE data (see Table 1). Furthermore, SHOT commonly outperforms model trained on dialectal data. This suggests that building dialect-robust models require the incorporation of SHOT-based self-supervised techniques in order to reverse the effect of performing self-supervised learning of base embedding models on SAE data.
Using Table 2, we calculate the individual dialectal gap and perform Pearson correlation analysis of these 27 observations (see Table 2) achieving a p-value of 0.0305 and a correlation coefficient of 0.4169. When considering only 9 observations with the train dataset being SAE, the p-value is 0.0041 and the correlation coefficient is 0.8455. In both cases, the p-value is statistically significant (<0.05) and the correlation coefficient is positive. The scatterplot is given in Figure 2.
Based on this evidence, we observe that there is a strong positive relationship between the dialectal gap and $\text{TTDA}_{\text{gain}}$. This means that as the dialectal gap increases, the more TTDA becomes beneficial in reducing said gap. In addition, When fine-tuning on a dialect $A$ (except for SAE) and evaluating on dialects $A,B_{1},B_{2},$..., the evaluation on dialect $A$ (the dialect the model was original fine-tuned on) has the best performance.
@@ 4 Conclusion
Due to the dialectal gap which exhibits in previous literature, we formally propose a mathematical definition of the concept, apply SHOT Liang et al. (2020) to dialectal NLP, and do a comparative study of SHOT against finetuning under different dialectal settings. Our findings indicate that:
  1. 1.
    When the model is fine-tuned on SAE as a source dialect, SHOT always improves the performance of models when adapting to the target dialect, even compared to models fine-tuned on the target dialect.
  2. 2.
    When the model is fine-tuned on a non-SAE source dialect $A$, it tends to have the highest performance when evaluating on the same target dialect $A$.
  3. 3.
    Dialectal gap has a strong positive relationship with $\text{TTDA}_{\text{gain}}$. This means that as the dialectal gap increases, the more TTDA is beneficial in reducing said gap.
Task Source Target Gap $\text{TTDA}_{\text{gain}}$
CoLA IndE NgE 1.260 3.16
CoLA IndE SingE 5.385 1.93
CoLA NgE IndE 3.220 3.85
CoLA NgE SingE 5.785 -0.47
CoLA SAE IndE 16.945 2.73
CoLA SAE NgE 14.965 2.02
CoLA SAE SingE 21.990 3.47
CoLA SingE IndE 0.000 2.73
CoLA SingE NgE -0.350 -1.26
RTE IndE NgE 0.800 -1.19
RTE IndE SingE 1.200 1.20
RTE NgE IndE 0.000 -3.18
RTE NgE SingE 1.190 -1.20
RTE SAE IndE -0.450 1.19
RTE SAE NgE -0.840 1.20
RTE SAE SingE -0.050 0.79
RTE SingE IndE 5.180 0.00
RTE SingE NgE 0.000 1.59
SST-2 IndE NgE 0.490 1.50
SST-2 IndE SingE 3.490 2.13
SST-2 NgE IndE 3.610 1.00
SST-2 NgE SingE 4.240 0.38
SST-2 SAE IndE 4.450 0.37
SST-2 SAE NgE 3.960 0.88
SST-2 SAE SingE 5.580 0.25
SST-2 SingE IndE 0.860 -0.88
SST-2 SingE NgE 0.860 0.74
Table 2: Dialectal gap and $\text{TTDA}_{\text{gain}}$. CoLA has a range of [1,1][-1,1] so we rescaled it to $[0,1]$. All gap values are in percentage. $\text{TTDA}_{\text{gain}}$ is between the TTDA and the baseline model as finetuned on SourceSource and evaluated on the TargetTarget.
Acknowledgments
The production of the code used in this paper was partially reliant on generative AI code assistant. Generative AI tools were used to perform literature search and for assisting in the wording of the paper. All AI-generation has been validated by the authors.
@@ Acknowledgments
Due to computational limits, we were not able to perform evaluation on all Multi-VALUE GLUE tasks. We also were not able to do experiments on STS-B due to this being a regression task which is not yet accommodated by SHOT Liang et al. (2020). Whilst computational constraints preclude tasks like QQP for the time being, we believe our chosen tasks sufficiently demonstrate the core contributions of our study (see Section 3.1). We used Multi-VALUE to create dialectal transformations which only produce syntactic variation and does not fully capture natural dialectal variation (see Section 2).
@@ Appendix A Appendix
T
IndE NgE SingE Average
S B A $\text{TTDA}_{\text{gain}}$ B A $\text{TTDA}_{\text{gain}}$ B A $\text{TTDA}_{\text{gain}}$ B A $\text{TTDA}_{\text{gain}}$
SAE 67.96 68.93 0.98 68.91 69.94 1.03 65.76 66.69 0.92 67.54 68.52 0.98
IndE 66.86 66.01 66.64 0.63 63.50 64.93 1.43 65.46 65.79 1.03
NgE 66.92 66.84 -0.09 69.20 65.46 65.11 -0.35 67.19 65.97 -0.22
SingE 66.34 66.50 0.16 68.18 68.75 0.57 68.35 67.62 67.62 0.36
Table 4: Average evaluation on CoLA, RTE, and SST-2. Bold is for best value for specific task and setting. Italic is for best performing evaluation dialect per dialect trained. S is the source dialect and T is the target dialect.
Figure 2: Scatterplot of the dialectal gap against $\text{TTDA}_{\text{gain}}$. The correlation coefficient is 0.4169 and the p-value is 0.0305. The scatter plot shows positive correlation between the dialectal gap and $\text{TTDA}_{\text{gain}}$, and this is truer for SAE-only observations for reasons pointed out above. We also plot regression line for SAE-only data and all data points